# SpiralShard: Highly Concurrent and Secure Blockchain Sharding via Linked Cross-shard Endorsement

### v1

@@ III-C Threat Model
In SpiralShard, there are honest and malicious nodes. The honest nodes obey all the protocols. There are two kinds of malicious nodes: Byzantine and alive-but-corrupt (a-b-c) nodes [36]. The Byzantine nodes may collude and corrupt the system arbitrarily, such as remaining silent, sending conflicting messages to different nodes, and voting for blocks containing invalid transactions. A-b-c nodes, proposed by Flexible BFT [36], are widely recognized as a typical and rational type of attacker, especially in real-world permissionless blockchains [27, 28, 15, 36]. A-b-c nodes always properly attempt to sign and send inconsistent messages to different nodes to break safety (e.g., fork chains to launch double-spending attacks). However, they positively vote for proposals to push consensus forward. Align with the previous work [36], a-b-c nodes should be treated as Byzantine during the safety proof. And once we have proved safety, we can treat a-b-c nodes as honest while proving liveness.
@@
A-b-c nodes are considered a typical and rational type of attacker [27, 28, 15, 36], especially in permissionless blockchain systems. The rationale behind this stems from the nature of blockchain payment systems, wherein some attackers often benefit by compromising system safety, evidenced by the substantial losses incurred due to double-spending attacks in Bitcoin and Ethereum Classic [3, 1]. On the other hand, nodes are motivated to preserve liveness to earn service fees [27, 28, 41, 15, 40, 42, 36].
The fraction of Byzantine and a-b-c nodes in the system are denoted as $F_{B}$ and $F_{A}$. Thus, there are $N\cdot F_{B}$ Byzantine and $N\cdot F_{A}$ a-b-c nodes in the system, respectively. Like most blockchain systems, we assume that the adversaries are slowly-adaptive, i.e., the set of malicious nodes and honest nodes are fixed during each epoch (e.g., one day) and can be changed only between epochs [34].
IV System design
@@
We aim to design a blockchain sharding system that achieves higher concurrency. As shown in Figure 2, we partition nodes in the network into multiple small shards, each maintaining its own shard chain. However, some shards may be corrupted due to their small size, compromising the system’s security (e.g., fork chains to launch double-spend attacks). Our intuition is to leverage honest shards to verify and endorse blocks on only one branch of forks in corrupted shards. Hence, we establish the concept of the endorsement group, which comprises multiple shards, including at least one honest shard, as shown in Figure 2. For the sake of safety and liveness, we have designed the LCE protocol that each shard sequentially declares its endorsement. As shown in Figure 2, the endorsements are linked to providing finality for ancestor blocks.
The running of SpiralShard proceeds in fixed periods called epochs. Like most existing sharding blockchain systems, we set the epoch duration to one day to resist the slowly-adapted adversary [17, 21, 46]. During each epoch, existing blockchain sharding systems perform intra-shard consensus within each shard to package transactions into blocks and securely append them to shard chains. However, since SpiralShard configures more shards to boost transaction concurrency, there may be corrupted shards in SpiralShard due to the smaller shard size. To tackle this issue, the main design difference between our system and previous systems is the Linked Cross-shard Endorsement protocol.
We first give the designs of the LCE protocol facing different challenges. Specifically, we illustrate the basic design of LCE preserving security in Section IV-A, how to optimize cross-shard communication for the LCE process in Section IV-B, and how to ensure the performance of the LCE protocol in Section IV-C. Subsequently, in Section IV-D, we introduce other system components, including the bootstrapping process, epoch transition, cross-shard transfer transactions, and cross-shard smart contracts. While these components are not the primary focus of our design innovations, they are integral parts of blockchain sharding systems.
@@
IV-A Basic Design of LCE
The LCE protocol requires that after a block is appended to its shard chain via intra-shard consensus, it must be sequentially endorsed by all shards in its endorsement group before being finalized. To illustrate the process of a block, from being appended to its shard chain to being finalized, the LCE protocol declares three states for each block: 1) prepared, 2) finalized, and 3) discarded. A block that passes the intra-shard consensus is considered prepared and would be finalized if it is endorsed by all shards from its endorsement group later. On the other hand, a prepared block that conflicts with finalized ones is considered discarded.
The basic design requires transmitting blocks across shards since they contain the history of each step of linked cross-shard endorsement. However, in our final design, only transmitting block headers is sufficient, reducing the cross-shard communication overhead. We will illustrate this later in Section IV-B. The procedure of LCE includes three phases:
Block Preparation. In SpiralShard, each shard must run BFT-typed intra-shard consensus to append blocks to its shard chain. We choose the Fast Byzantine Fault Tolerance (FBFT[5] consensus protocol proposed by Harmony as our intra-shard consensus protocol. FBFT is a variant of Practical Byzantine Fault Tolerance (PBFT[12], adopting a short signature scheme, Boneh-Lynn-Shacham (BLS) multi-signature [10] to collect votes from participants for higher scalability. Besides, similar to PBFT [12], FBFT has a review change mechanism that relies on $2/3$ of nodes’ participation to replace the Byzantine leader of consensus. The intra-shard consensus protocol is a pluggable component as long as it uses a $2/3$ quorum size and tolerates less than $1/3$ Byzantine nodes (e.g., PBFT [12]). Therefore, nodes can verify whether a block has passed the intra-shard consensus through the signatures in the block header. Such a block that passes the intra-shard consensus is valid (does not conflict with the parent block) and considered prepared and receives the endorsement from its shard. Honest nodes must broadcast the prepared block they vote for to the next shard for cross-shard endorsement. An honest shard always appends only one prepared block to the tail of its shard chain each time. However, when a shard is corrupted, it can append multiple prepared blocks after the same block, resulting in a fork.
@@ IV-A Basic Design of LCE
Cross-shard Endorsement. The LCE protocol utilizes the endorsement group of multiple shards to decide on one of the fork branches on a corrupted shard. Since an honest shard will only endorse blocks on one fork branch on a corrupted shard to eliminate forks, a group must contain at least one honest shard with high probability. To ensure that multiple honest shards within the same endorsement group make consistent endorsement decisions, the LCE protocol requires that shards conduct endorsements based on the ascending order of Shard IDs.
@@
For example, as shown in Figure 3, without loss of generality, an endorsement group consists of $Shard_{0}$, $Shard_{1}$, and $Shard_{2}$. After blocks $A_{2}^{\prime}$ and $A_{2}^{\prime\prime}$ are prepared, $Shard_{1}$ and $Shard_{2}$ need to determine their endorsements in succession. In $Shard_{1}$, block $B_{3}$ can endorse either block $A_{2}^{\prime}$ or $A_{2}^{\prime\prime}$ without conflict because block $B_{2}$ (parent of block $B_{3}$) has endorsed block $A_{1}$ (parent of block $A_{2}^{\prime}$ and $A_{2}^{\prime\prime}$). Without loss of generality, the proposer of block $B_{3}$ in $Shard_{1}$ receives valid prepared block $A_{2}^{\prime}$ first, so $Shard_{1}$ endorses block $A_{2}^{\prime}$ in block $B_{3}$. In corrupted $Shard_{2}$, the malicious block proposer forks the shard chain by proposing two blocks ($C_{4}^{\prime}$ and $C_{4}^{\prime\prime}$), both of which can only endorse $B_{3}$ after $Shard_{1}$’s endorsement decision.
@@
IV-B Efficient Cross-shard Communication for LCE

### v2

@@ III-B Threat Model
In SpiralShard, there are honest and malicious nodes. The honest nodes obey all the protocols. There are two kinds of malicious nodes: Byzantine and alive-but-corrupt (a-b-c) nodes [flexible]. The Byzantine nodes may collude together (both intra-/cross-shard) and corrupt the system arbitrarily, such as remaining silent, sending conflicting messages to different nodes, and voting for blocks containing invalid transactions. Besides, bounded by the eventual synchrony property of the partial-synchronous network, Byzantine nodes can only exploit short-term message delays or reorderings. A-b-c nodes, proposed by Flexible BFT[flexible] (CCS ’19), are widely recognized as a typical and rational type of attacker, especially in real-world permissionless blockchains [jo2020toward, kane2021highway, Trebiz, flexible]. A-b-c nodes always properly attempt to sign and send inconsistent messages to different nodes to break safety (e.g., generate fork chains). However, they positively vote for proposals to push consensus forward (i.e., maintain liveness). Align with the previous work [flexible], a-b-c nodes should be treated as Byzantine during the safety proof. And once we have proved safety, we can treat a-b-c nodes as honest while proving liveness.
@@
Justification of A-b-c Nodes. A-b-c nodes are considered a typical and rational type of attacker [jo2020toward, kane2021highway, Trebiz, flexible], especially in permissionless blockchain systems. The rationale behind this stems from the nature of blockchain payment systems, wherein some attackers often benefit by compromising system safety, evidenced by the substantial losses incurred due to double-spending attacks in Bitcoin and Ethereum Classic [etcdouble]. On the other hand, nodes are motivated to preserve liveness to earn service fees [jo2020toward, kane2021highway, ranchal2020zlb, Trebiz, blockchainIsDead, basilic, flexible].
SpiralShard is a blockchain sharding system that employs a BFT-typed consensus mechanism within each shard. Operating under a partially synchronous network model, SpiralShard has an overall fault tolerance threshold of 1/3. Specifically, to maintain system security, the total fraction of malicious nodes $F$ (including Byzantine and a-b-c nodes) across the entire network must remain below 1/3. In other words, the total number of malicious nodes are $N\cdot F$. The fraction of Byzantine and a-b-c nodes in the system are denoted as $F_{B}$ and $F_{A}$, and $F_{B}+F_{A}=F$. Correspondingly, there are $N\cdot F_{B}$ Byzantine and $N\cdot F_{A}$ a-b-c nodes in the system, respectively.
Slowly-Adaptive Adversaries. Like most blockchain systems [rapidchain, omniledger, li2022jenga, repchain, sgxSharding, pyramid, brokerchain], the adversaries in the system are slowly-adaptive, i.e., the set of malicious nodes and honest nodes in each shard are fixed during each epoch and can only be changed after shard reconfiguration [buildingBlocks_survey] (see Section IV-C). In other words, malicious nodes may attempt to corrupt honest nodes (e.g., through bribery or other means). However, such adaptive attacks require at least one epoch to take effect. Due to the randomized shard reconfiguration process conducted at each epoch transition, these slowly adaptive malicious nodes cannot compromise this randomness, thereby preventing the aggregation of an excessive number of malicious nodes within any particular shard.
IV System Overview and Objectives
IV-A Overview of Main Designs
Existing large-scale permissionless blockchain sharding systems typically require each shard to be honest (e.g., maintaining <1/3<1/3 fraction of malicious nodes within each shard for BFT-type consensus under partially synchronous networks) to guarantee the security of the entire system. However, from a probabilistic perspective, due to the random distribution of nodes into shards, ensuring that every shard remains honest except with negligible failure probability necessitates configuring excessively large shard sizes. Such oversized shard configurations inevitably reduce the transaction concurrency of the system.
To enhance system performance, SpiralShard first allows the existence of some corrupted shards (where $\geq 1/3$ fraction of malicious nodes within a shard). Consequently, under the same network size and overall proportion of malicious nodes, it becomes feasible to configure more and smaller shards, thus significantly improving transaction concurrency. Second and more importantly, to simultaneously boost transaction concurrency and maintain security, we innovatively propose the Linked Cross-shard Endorsement (LCE) protocol. LCE acts as a secondary security protocol operating across shards on top of intra-shard consensus. Its core principle is leveraging the collaborative effort of multiple shards to recover corrupted shards and safeguard overall system security.
Specifically, as illustrated in Figure 2, multiple shards, including at least one honest shard, constitute an endorsement group, collectively securing each shard within the group. To ensure (except with negligible failure probability) that each endorsement group contains at least one honest shard, we calculate appropriate shard sizes and endorsement group sizes via rigorous probabilistic modeling and mathematical derivation (Section VI-A). Multiple such endorsement groups coexist within the system simultaneously. Shards within an endorsement group individually run their intra-shard consensus and sequentially performs cross-shard endorsements on the endorsement results from its preceding shard, strictly adhering to a predetermined order. Consequently, any intra-shard consensus outcome obtains sequentially linked endorsements. Once a shard’s consensus outcome has been sequentially and fully endorsed by all shards in its endorsement group, it can be considered finalized and secure (Section V-A). This is primarily because the presence of the honest shard within the endorsement group ensures it will neither engage in malicious behavior nor endorse fraudulent endorsements.
Furthermore, SpiralShard employs a mixed threat model and precisely adjusts the shard sizes (again, through rigorous probabilistic derivations (Section VI-A)) to ensure, except with negligible failure probability, that each shard’s fraction of malicious nodes remains below 2/3 (as opposed to the conventional threshold of 1/3). This includes less than 1/3 Byzantine nodes and up to 1/3 alive-but-corrupt (a-b-c) nodes. This approach provides two primary benefits. First, the mixed threat model enhances system liveness because a-b-c nodes solely target safety, not liveness. Second and more importantly, limiting the proportion of malicious nodes within each shard to below 2/3 significantly reduces the overhead associated with cross-shard endorsement. At this threshold, adversaries cannot control enough malicious nodes within a shard to let invalid transactions pass intra-shard consensus, thus ensuring transaction validity internally within each shard. Nonetheless, adversaries may still execute equivocation attacks to cause forks. Fortunately, such forks can be efficiently detected using lightweight block headers. Consequently, the LCE protocol enables the transmission of only block headers (rather than entire blocks) as intra-shard consensus results for cross-shard communication. As a result, the LCE protocol can effectively resolve intra-shard forks through connected cross-shard endorsements at minimal overhead, thereby securing the overall system (Section V-B).
@@ IV-A Overview of Main Designs
IV-B Objectives
SpiralShard aims to achieve the following objectives:
  • Security: Given that the total fraction of malicious nodes $F$ within the entire system is below its fault tolerance threshold (see Section III-B), SpiralShard aims to guarantee system security (i.e., safety and liveness) except with negligible failure probability. In this paper, we define the negligible failure probability as $2^{-17}$, indicating the system is expected to fail only once every 359 years for one-day epoch. This definition is consistent with many existing blockchain sharding systems [pyramid, li2022jenga, rapidchain, omniledger, cochain]. Specifically, for each shard, the system ensures—except with negligible failure probability—that it contains fewer than 2/3 malicious nodes, including fewer than 1/3 Byzantine nodes and at most 1/3 a-b-c nodes. For each endorsement group, the design ensures—except with negligible failure probability—the inclusion of at least one honest shard (i.e., a shard with a malicious node proportion below 1/3). After achieving these two objectives, the Linked Cross-shard Endorsement (LCE) protocol further aims to guarantee both safety (i.e., honest nodes agree on the sequence of the same and valid finalized blocks in each shard chain [rapidchain, omniledger]) and liveness (i.e., transactions in each shard are finalized within finite time, the system does not stall indefinitely [rapidchain, omniledger]).
  • High Concurrency: By securely allowing shards to contain a higher fraction of malicious nodes (<2/3<2/3 rather than the conventional <1/3<1/3), SpiralShard aims to configure a greater number of smaller-sized shards within the same network scale and overall malicious node ratio. On one hand, this increases the total number of shards in the system; on the other, it enhances the efficiency of intra-shard consensus. Through these combined effects, SpiralShard aims to significantly boost overall system throughput (transactions per second, TPS), achieving higher transaction concurrency.
The proofs, analysis, and experiments related to the aforementioned objectives are detailed in Section VI and Section VII.
IV-C Overview of Other Components
Beyond the overview of our main designs, we now provide brief remarks about other notable components in the system.
Epochs. SpiralShard operates based on epochs of fixed duration. The exact length of an epoch may vary according to specific system implementations, but without loss of generality, we set the epoch length to one day, consistent with many existing systems [sgxSharding, pyramid, rapidchain]. Within each epoch, every shard continuously runs its intra-shard consensus to advance the system’s state. More importantly, each shard in SpiralShard simultaneously executes the LCE protocol to ensure system security.
Shard Reconfiguration. Between epochs, SpiralShard performs standard shard reconfiguration (consistent with most existing blockchain sharding systems [rapidchain, pyramid, cycledger, Harmony]) to reassign nodes among shards. This procedure primarily aims to defend against slowly-adaptive adversaries, preventing them from gradually corrupting individual shards. For security, the nodes are randomly assigned to different shards. Specifically, a random seed is generated at the end of each epoch (e.g., by a decentralized randomness protocol or via the beacon chain, like RANDAO in Ethereum 2.0 or other ways [eth, Harmony]). Similar to many existing sharding systems [omniledger, rapidchain, pyramid, li2022jenga, Harmony], using this seed, a VRF (Verifiable Random Function) [micali1999verifiable] is run by each node to determine if it is selected to randomly move to a different shard (or to remain). Additionally, a VDF (Verifiable Delay Function) [boneh2018verifiable] can be used to make the randomness unpredictable until it is time to apply it, to prevent adversaries from precomputing favorable placements. In this way, we can guarantee that the node reassignment of each shard during the shard reconfiguration is random, verifiable, unbiased, and unpredictable. Due to this randomness, the adversary cannot ensure their nodes stay together in one shard after reconfiguration.
Cross-shard Transactions. SpiralShard can leverage existing mechanisms to handle cross-shard transactions, even when these transactions span across different endorsement groups. However, in SpiralShard, to ensure the security of cross-shard transactions, we require that cross-shard transactions be processed and finalized in other shards only after they are finalized by the LCE protocol. Specifically, the cross-shard transfer transactions are processed via the existing relay-based mechanism [Harmony, monoxide]. Moreover, our system is inherently capable of supporting smart contracts, primarily because we ensure the safe finalization of blocks. To facilitate cross-shard contract transactions, we can leverage the traditional Two-Phase Commit (2PC) protocol [sgxSharding, li2022jenga]. The detailed discussion about the cross-shard smart contract processing and its limitation are given in Section VI-D.
V System design
We first introduce the basic LCE protocol in Section V-A. Building upon this, Section V-B describes how we leverage block headers to reduce the overhead associated with cross-shard endorsements. Finally, Section V-C presents a pipelining mechanism designed to safely and further enhance the protocol’s performance.
@@ V System design
V-A Basic Design of LCE
The LCE protocol requires that after a block is appended to its shard chain via intra-shard consensus, it must be sequentially endorsed by all shards (via each shard’s intra-shard consensus) in its endorsement group before being finalized. To illustrate the process of a block, from being appended to its shard chain to being finalized, the LCE protocol declares three states for each block: 1) prepared, 2) finalized, and 3) discarded. A block that passes the intra-shard consensus is considered prepared and would be finalized if it is endorsed by all shards from its endorsement group later. On the other hand, a prepared block that conflicts with finalized ones is considered discarded.
The basic design requires transmitting blocks across shards since they contain the history of each step of linked cross-shard endorsement. However, in our final design, only transmitting block headers is sufficient, reducing the cross-shard communication overhead. We will illustrate this later in Section V-B. The procedure of LCE includes three phases:
Block Preparation. In SpiralShard, each shard runs BFT-typed intra-shard consensus to append blocks to its shard chain. During the consensus, nodes can verify whether a block has passed the intra-shard consensus through the signatures in the block header. Such a block that passes the intra-shard consensus is valid (does not conflict with the parent block) and considered prepared and receives the endorsement from its shard. Honest nodes must broadcast the prepared block they vote for to the next shard for cross-shard endorsement. An honest shard always appends only one prepared block to the tail of its shard chain each time. However, when a shard is corrupted, it can append multiple prepared blocks after the same block, resulting in a fork.
@@ V-A Basic Design of LCE
Choice of Intra-shard Consensus. The intra-shard consensus protocol is a pluggable component in SpiralShard, as long as it is a standard BFT-typed consensus and can operate in a partial-synchronous network (it uses a $2/3$ quorum size and tolerates less than $1/3$ Byzantine nodes, like Practical Byzantine Fault Tolerance (PBFT[castro1999practical]). In this work, we choose the Fast Byzantine Fault Tolerance (FBFT[Harmony] consensus protocol proposed by Harmony [Harmony] as our intra-shard consensus protocol. FBFT is a variant of PBFT [castro1999practical], adopting a short signature scheme, Boneh-Lynn-Shacham (BLS) multi-signature [bls] to collect votes from participants for higher scalability.
FBFT has a view change mechanism to replace the Byzantine leader of consensus, similar to PBFT [castro1999practical]. Moreover, like existing BFT-type consensus protocols in partially synchronous networks, FBFT also incorporates a timeout mechanism to trigger the view change process for liveness. Similarly, during the asynchronous phase before GST, the timeout duration gradually increases with each view change and eventually converges to a stable value (once the network becomes synchronous), to accommodate the asynchronous nature of the network and prevent overly frequent view changes.
After GST, the timeout in SpiralShard converges to a small constant value (on the order of only a few seconds). This behavior is consistent with standard partially synchronous consensus protocols like PBFT and Tendermint [castro1999practical, buchman2016tendermint], and we inherit it in SpiralShard’s implementation (using FBFT [Harmony]) to ensure timely view changes with minimal cost. As a result, any leader replacement (view change) after GST completes rapidly, incurring minimal overhead. Additionally, under normal conditions, FBFT regularly rotates leaders (e.g., once per epoch) based on the identity of each consensus node, aligning with many existing BFT-type consensus protocols [buchman2016tendermint, Harmony]. Since we adopt a leader rotation approach similar to existing BFT-type consensus mechanisms and further ensure the overall system’s security through the LCE protocol, the leader replacement process in SpiralShard is also secure. A detailed proof is presented in Section VI-B.
Cross-shard Endorsement. The LCE protocol utilizes the endorsement group of multiple shards to decide on one of the fork branches on a corrupted shard. Since an honest shard will only endorse blocks on one fork branch on a corrupted shard to eliminate forks, a group must contain at least one honest shard except with negligible failure probability. To ensure that multiple honest shards within the same endorsement group make consistent endorsement decisions, the LCE protocol requires that shards conduct endorsements based on the ascending order of Shard IDs.
@@
Fork Selection. In the LCE protocol, the choice of which fork to endorse is crucial, as it determines which branch within a corrupted shard will ultimately be finalized by the LCE protocol. In a partially synchronous network, deterministically identifying the correct fork is impossible. Fortunately, SpiralShard does not require deterministic fork detection. In the LCE protocol, the leader follows two main rules when selecting a fork from the previous shard. First, the selected fork must be an extension of the previously finalized fork as determined by the LCE protocol, to prevent conflicts. Second, there may be multiple forks that extend from the same finalized fork. Among these forks, the leader will select the longest one from its own perspective, and then propose the latest block on that fork for endorsement along with other transactions through intra-shard consensus. For the honest nodes in the shard, they verify during consensus whether the fork chosen by the leader is indeed an extension of the previously finalized fork and whether the leader selected only one fork for endorsement (to prevent equivocation attacks). However, the honest nodes do not verify if the leader has chosen the longest fork, since the information observed by different nodes may be uncertain in a partially synchronous network. This fork selection mechanism avoids the limitations of nondeterminism in a partially synchronous network while ensuring the security of the protocol.
Security of Fork Selection. During the fork selection process, a malicious leader may attempt to violate the fork selection rules through various means. However, these attempts do not compromise the security of the LCE protocol. For instance, a malicious leader in an honest shard may attempt to select multiple forks (equivocation attack), choose a conflicting fork (i.e., one that is not an extension of the previously finalized fork by the LCE protocol), or refuse to produce a block. However, as mentioned, since all endorsements must be approved through intra-shard consensus, such malicious actions will be detected by honest nodes within the shard, resulting in endorsement failure and triggering a view change. A malicious leader in an honest shard may also attempt to avoid endorsing the longest fork from its perspective. This behavior can be tolerated, as long as the leader attempts to endorse a block from any non-conflicting fork (and not multiple forks), the honest nodes in intra-shard consensus do not determine whether the fork is the longest. Therefore, this selection will pass consensus and will not compromise the protocol’s security, although it may slow down the system’s progress. For corrupted shards, consensus within the shard may be used to endorse multiple forks against the rules. However, since there is at least one honest shard within each endorsement group, as described earlier, that honest shard ensures that a single fork is selected for endorsement through intra-shard consensus, thereby ensuring the overall security of the LCE protocol. A simple example is illustrated in Figure 3.
As shown in Figure 3, without loss of generality, an endorsement group consists of $Shard_{0}$, $Shard_{1}$, and $Shard_{2}$. After blocks $A_{2}^{\prime}$ and $A_{2}^{\prime\prime}$ are prepared, $Shard_{1}$ and $Shard_{2}$ need to determine their endorsements in succession. In $Shard_{1}$, block $B_{3}$ can endorse either block $A_{2}^{\prime}$ or $A_{2}^{\prime\prime}$ without conflict because block $B_{2}$ (parent of block $B_{3}$) has endorsed block $A_{1}$ (parent of block $A_{2}^{\prime}$ and $A_{2}^{\prime\prime}$). Without loss of generality, the proposer of block $B_{3}$ in $Shard_{1}$ receives valid prepared block $A_{2}^{\prime}$ first, so $Shard_{1}$ endorses block $A_{2}^{\prime}$ in block $B_{3}$. In corrupted $Shard_{2}$, the malicious block proposer forks the shard chain by proposing two blocks ($C_{4}^{\prime}$ and $C_{4}^{\prime\prime}$), both of which can only endorse $B_{3}$ after $Shard_{1}$’s endorsement decision.
@@
V-B Efficient Cross-shard Communication for LCE